OpenStudy (anonymous):

can anyone show me how to work x^6=46656 ?

OpenStudy (anonymous):

Factor it.

OpenStudy (anonymous):

Keep dividing by 2. Then keep dividing by 3. It will be clear once you factor it.

OpenStudy (anonymous):

You will end up with \[ 46656 =2^6\times 3^6 \]

OpenStudy (anonymous):

is that the final answer?

OpenStudy (anonymous):

No, but that should make it pretty clear what \(x\) is.

OpenStudy (anonymous):

6 ?

OpenStudy (anonymous):

You didn't ask for an answer, you asked for how to work it.

OpenStudy (anonymous):

Yes, \(x=6\)

OpenStudy (anonymous):

yea i just wanted to make sure I was right thank you

OpenStudy (wolf1728):

How about taking logarithms? x^6=46656 log x^6 = log 46656 6 * log x = log 46656 I'm using base 10 logs here: log x = log 46656 / 6 log x = 4.668907502302 / 6 log x = 0.7781512504 10^0.7781512504 = x x = 6